OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{x}{y^2\sqrt{1+x}}\]

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{x}{y^2\sqrt{1+x}}\]\[y^2dy = \frac{x}{\sqrt{1+x}}dx\]\[\frac{1}{3}y^3= \int \frac{u-1}{u}du = \frac{2}{3} u^{\frac{3}{2}}- 2u^{\frac{1}{2}}+C\]\[y = 2(1+x)^{\frac{3}{2}}-6(1+x)^{\frac{1}{2}}+C)^{\frac{1}{3}}\]

OpenStudy (anonymous):

Did I do something wrong again?

hartnn (hartnn):

i couldn't understand clearly what u did, but it can be simplified to 2/3*(x-2)*sqrt(x+1)

OpenStudy (anonymous):

Which part you didn't understand?

hartnn (hartnn):

u=?

OpenStudy (anonymous):

1+x

hartnn (hartnn):

then denominator is sqrt u

OpenStudy (anonymous):

OMG!I typed something wrong!! \[\frac{1}{3}y^3= \int \frac{u-1}{\sqrt u}du = \frac{2}{3} u^{\frac{3}{2}}- 2u^{\frac{1}{2}}+C\]

OpenStudy (anonymous):

No wonder why you couldn't understand.. That was my mistake :( Is that clear now?

hartnn (hartnn):

yeah...u can simplify your final answer by factoring out sqrt(x+1)

hartnn (hartnn):

if needed.

OpenStudy (anonymous):

That's how the answer in my book looks like :S But thanks!

hartnn (hartnn):

ohh.. then no need. you were correct all along. welcome ^_^